# Detection and Tracking of MAVs Using a Rosette Scanning Pattern LiDAR

### v2

@@ III-E Implementation details
The Robot Operating System (ROS) was used to integrate and manage communication between program components, offering a reliable framework for modular nodes. Three nodes were implemented: a LiDAR wrapper node to fix point cloud conversion issues, a pan-tilt turret wrapper node to translate MAV position estimates into turret commands, and a MAV detection and tracking node that holds the core algorithm, including the particle filter.
IV INDOOR EXPERIMENTS and EVALUATION
First, the experimental results are presented and examined, starting with the initial detection of the MAV following the construction of the background model. Next, the core of the algorithm, the tracking process, is evaluated in terms of accuracy. Finally, the re-detection capability after tracking loss is assessed.
The experiments were conducted within the MIMO arena [26] within the HUN-REN SZTAKI Institute for Computer Science and Control incorporating an OptiTrack motion capture system in a $7\times 8\;m^{2}$ office space with a height of $3.5\;m$. Further hardware includes MAVs (Crazyflie 2.0 by Bitcraze), the turret (WidowX Dual XM430 Pan & Tilt turret by Trossen Robotics), and the rosette scanning pattern LiDAR (Avia LiDAR sensor by Livox) with an integration time of $100\;ms$ ($10\;Hz$).
In all our experiments the particle filter was run at $15\;Hz$ frequency which assured the processing of all LiDAR measurements. The processing ran on an 8th-generation desktop Intel i7 processor with less than 30% utilization.
IV-A Detecting the MAV
@@
Figure 2: Initial detection error and $\sigma_{particles}$ of the particle filter. The particle filter locks onto the drone almost instantaneously after receiving the first measurement at around 2 seconds.
@@
Fig. 2 shows the error as the distance of the estimated state to the true state in meters of a representative detection. The average of the standard deviation of the particles in each dimension is also plotted (denoted as $\sigma_{particles}$).
When there is no incoming measurement only the predict step of the particle filter is performed. In all of our experiments $\sigma_{pred}=0.1\;m$ and $n=500$, which turned out to be a good compromise between being able to track fast-moving drones and providing accurate measurements.
@@
While the drone is out of the FoV of the LiDAR sensor, the particles remain idle, only influenced by the added noise in the predict step which results in a slowly inflating particle cloud with a slowly rising $\sigma_{particles}$. In this case, the MAV is resting on the floor and, because of its small size, is hidden inside the background model. As the drone starts hovering and measurements of it arrive at around 2 s the particles gather and allow for the first good estimation. Parallel with this the $\sigma_{particles}$ of the estimation drops sharply. This means the drone is tracked and the particle filter also predicts velocity.
IV-B Tracking the MAV
@@
For calculating the accuracy, the MAV was flown following certain flight patterns (described in Fig. 3) in a repeated manner. The ground truth data was obtained by the motion capture system within the MIMO arena by recording the pose of the LiVOX sensor and the position of the drone and aligning the LiVOX measurements with the MIMO arena using the recorded pose. The patterns in Fig. 3 were drawn from the point of view of the LiDAR sensor where the middle of each trajectory in the horizontal direction was in front of the sensor with a minimum distance of around $3\;m$. Each circle represents a stationary wait time of $2\;s$ and the one with a black outline represents the first position of the respective pattern. The trajectory represented by the arrows in-between was flown in $3\;s$ in case of the vertical, horizontal, and ’lost and found’ pattern which resulted in a maximum speed of about $0.6\;m/s$, while it was flown in $2.25\;s$ in case of the fast pattern which resulted in a maximum speed of around $1.2\;m/s$. In each experiment, the drone flies to the first point and then it flies the pattern 3 times in the case of the vertical and horizontal patterns and 4 times in the case of the others.
@@
In Fig. 4, the error of the particle filter is plotted with the ground truth speed calculated from the MIMO measurements. It is visible that the tracking is stable over time with an average position error of under $5\;cm$ in the case of the horizontal and vertical measurements while it is around $7\;cm$ for the fast pattern.
@@
Figure 4: Position error of the particle filter plotted together with ground truth speed from the MIMO arena for the vertical (top), horizontal (middle), and fast (bottom) patterns. The mean ($\mu_{error}$) and standard deviation (±$\pm\sigma_{error}$) of the particle filter error are also plotted.
@@
It can be seen on all the plots of Fig. 4, that the error of the particle filter is dependent on the speed. When the drone is at a waypoint, not moving, the error of the filter is around $3$ to $5\;cm$, while at the peak speeds of $1.2\;m/s$ of the fast pattern (not counting the one outlier), the error is around $20\;cm$. This dynamic difference of about $15\;cm$ in the error can mostly be explained by the different delays in the pipeline. The internal processing time of the LiDAR firmware is $2\;ms$ based on the data-sheet, the LiDAR scan integration time is $100\;ms$, and there is some additional overhead in the formatting, stamping, and publishing of the data to ROS topics. The background subtraction and the application of filters took $10.9\;ms$ with a $\sigma=16.8\;ms$ standard deviation on the fast trajectory and the predict, update and resample steps took a combined red1.$1.11\;ms$ with a $\sigma=1.12\;ms$. The whole delay is about $120\;ms$, in which the drone travels roughly $14.4\;cm$. From the elapsed time, the computation takes only around $20\;ms$ resulting in convenient real-time operation.
The particle filter’s inherent functionality causes the particles to consistently lag slightly behind the measurements. Additionally, the relative motion between the drone and the LiDAR sensor naturally introduces smearing in the point cloud, accounting for the remaining dynamic error.
@@
Overall, when the drone is stationary, the error remains minimal and consistent. However, during movement, a dynamic error arises that increases linearly with speed. This affects the tracking accuracy but does not compromise its robustness.
For comparison, in [24] the RMSE is reported as the metric of accuracy on different trajectories. From their paper, the trajectory T1 is most similar to our horizontal pattern and the trajectory T3 is most similar to our vertical case. We consider them similar because the trajectories are mostly in the horizontal and vertical planes respectively and mostly contain straight lines and turns close to right angles. Their reported best RMSE for T1 and T3 are $0.0773\;m$ and $0.0421\;m$ while the RMSE for our horizontal and vertical patterns are $0.0570\;m$ and $0.0531\;m$ respectively. While the two methods work very similarly in terms of accuracy, our particle filter approach has the added benefit of being able to regain the tracking after it was lost.
IV-C Losing and regaining Tracking

### v3

@@ II-E Implementation details
The Robot Operating System (ROS) was used to integrate and manage communication between program components, offering a reliable framework for modular nodes. Four nodes were implemented: a LiDAR wrapper node to fix point cloud conversion issues; a pan-tilt turret wrapper node to translate MAV position estimates into turret commands and send them to the turret’s position controller; a background builder node for the Octomap background model; and a MAV detection and tracking node that holds the core algorithm, including the particle filter and other Extended Kalman Filter (EKF) implementations for benchmarking. Note that the source code is made available8.
III Experiments
In the following subsections, we explain the parameter settings and discuss the experimental results, including a comparison with other methods. We first present and validate the equations for setting the parameters of the different motion models of the particle filter in Subsection III-A. In Subsection III-B, we quantify the effect of the different parts of our algorithm. Then, in Subsection III-C, an extensive comparison is carried out on four recordings, comparing our algorithm with other SOTA methods [13, 24]. A real-world field test is presented in Subsection III-D. The accuracy of the proposed method is investigated in III-E. Finally, we show the track regain capability of our algorithm in III-F.
In all of the experiments, the turret used was the WidowX Dual XM430 Pan & Tilt turret by Trossen Robotics, and the integration time of the LIVOX AVIA sensor was set to 100ms100\;ms.
III-A Motion model parameter selection
The particle filter relies on several key parameters: the number of particles $n$ and noise parameters $\sigma_{pos}$, $\sigma_{vel}$, $\sigma_{acc}$, $\sigma_{m}$, which govern how particles evolve and interact with the measurements. We justify our parameter choices based on computational constraints, drone dynamics, and experimental validation.
To ensure real-time operation, we evaluated the effect of particle count $n$ on computational load using the Tracy Profiler [^9] https://github.com/wolfpld/tracy. The filter was benchmarked on a single thread of an Intel Core i7 (8th generation) desktop CPU. The LiDAR callback can take up a worst-case execution time of 52ms52\;ms (but usually takes much less: 1.45ms1.45\;ms mean with 1.57ms1.57\;ms standard deviation), regardless of the number of particles so each filter step (prediction, update, and resampling combined) had to complete within $48ms$ to maintain a 10Hz10\;Hz filter rate that is needed to process the LiDAR data at 10Hz10\;Hz. TABLE I presents the mean, standard deviation, and worst-case execution times of the particle filter across different particle counts and the worst-case full execution time with the LiDAR callback, which has to be under 100ms100\;ms to be able to run at 10Hz10\;Hz.
TABLE I: Execution times with different particle numbers. Worst-case values are significant outliers, as indicated by the mean and standard deviation.
n [][-] 10000 9000 8000 7000 6000 5000 4000
Mean PF $[ms]$ 20.60 17.30 16.03 15.38 12.5 10.85 10.55
Std PF $[ms]$ 2.52 3.56 1.88 4.54 3.19 3.13 4.81
Worst-case
PF $[ms]$
103.00 88.32 81.38 59.11 50.67 34.62 33.16
Worst-case
full $[ms]$
155.00 140.32 133.38 111.11 102.67 86.62 85.16
The particles approximate the probability distribution of the model. The accuracy of this approximation improves with increasing particle count, asymptotically approaching the true distribution. The number was maximized while still enabling real-time operations. Based on these results, we set $n=5000$ to ensure real-time performance while maintaining a sufficiently accurate representation of the posterior distribution.
For context, we report the EKF execution times of the two SOTA methods. The mean, standard deviation, and worst-case execution times are (0.12[ms]; 0.11[ms]; 1.73[ms])(0.12\;[ms];\,0.11\;[ms];\,1.73\;[ms]) for [13] and (0.12[ms]; 0.16[ms]; 3.03[ms])(0.12\;[ms];\,0.16\;[ms];\,3.03\;[ms]) for [24]. While these results indicate noticeably lower execution times, the difference is not significant in the context of the overall pipeline and is justified by the performance improvements outlined in the following Subsections.
We tested all three motion models for particle prediction. Each model predicts particles based on different assumptions and therefore requires different parameter sets. Initial tests revealed that while the delta position and constant velocity models performed comparably, the constant acceleration model consistently underperformed even with small $\sigma_{acc}$ values. This was due to over-inflation of the particle cloud and poor handling of abrupt directional changes, issues caused by the assumption of continuous acceleration. As a result, it was not considered for further evaluation.
Based on the results of these initial experiments, we derived a unified set of equations for initializing the parameters of all models, grounded in the expected dynamics of the drone. This unification promotes consistency and simplifies the tuning process.
The equations below define $\sigma_{pos}$, $\sigma_{vel}$, $\sigma_{acc}$, and $\sigma_{m}$ in terms of known or measurable drone dynamics, namely expected maximum speed $v_{max}$ and expected maximum acceleration $a_{max}$.
$\sigma_{pos}$ $=\frac{v_{max}\cdot T}{3}$ (11)
$\sigma_{vel}$ $=\frac{a_{max}\cdot T}{3}$
$\sigma_{acc}$ $=0$
$\sigma_{m}$ $=\frac{s_{diag}}{2}$
In (11), $T$ is the filter’s time step and $s_{diag}$ is the diagonal distance between the drone’s motors. The choice of dividing by 3 ensures that approximately 99.74% of the particles remain within the physically plausible range (assuming a Gaussian distribution). The $\sigma_{acc}$ parameter is set to 0 because there is no constantly changing acceleration in the drone’s dynamics. Finally, $\sigma_{m}$ was chosen similarly to [24].
To validate these parameters, we conducted flight experiments with a DJI Phantom 4 ($s_{diag}=0.4\;m$). The drone was flown along a challenging trajectory, including high-speed passes perpendicular to the LiDAR’s axis and out-of-range segments to test edge cases that can be seen in FIGURE 2 on the far right. Ground Truth (GT) positions were obtained via GPS, streamed in real-time to the processing laptop using UGCS [^10] https://www.ugcs.com/, where the turret was moved based on these GT measurements. Despite minor latency, the setup provided reliable ground truth for evaluation.
Figure 2: Returned point counts and trajectories are shown for the four different evaluation flights. The color indicates the number of MAV points in the point cloud from the LIVOX sensor, with red representing zero returns. The ’Simple distance’, ’Horizontal slow’, and ’Horizontal fast’ trajectories were flown at constant altitude and are therefore plotted from a top-down perspective. The ’Vertical’ trajectory was flown in a line varying the altitude, and is plotted from a side view. All trajectories were designed such that, at the maximum distance, no points are returned for extended periods. The maximum detection distance is indicated above each figure, with a small black arrow also marking this point along the trajectory.
The maximum observed velocity and acceleration were approximately 15m/s15\;m/s and $15\;m/s^{2}$, leading to the following parameters based on (11): $\sigma_{pos}=0.5\;m$, $\sigma_{vel}=0.5\;m/s$, $\sigma_{acc}=0\;m/s^{2}$, and $\sigma_{m}=0.2\;m$. This speed and acceleration represent the upper limit at which our pilot was still comfortable controlling the drone, and is therefore considered a good approximation of realistic conditions in practical scenarios.
We then performed a grid search to evaluate the system’s sensitivity to parameter variations:
[…]
Each parameter combination was evaluated across 5 trials to account for the non-deterministic nature of the particle filter. We used two metrics: (1) maximum tracking distance, defined as the average distance of the furthest correctly identified MAV point cloud (within 4m4\;m of GPS ground truth), and (2) track coverage, the percentage of the flight where the MAV was correctly tracked.
FIGURE 3 shows a subset of the results for the constant acceleration model for illustration purposes. Points are color-coded by the Euclidean distance from the analytical parameter set given by equations (11). Results indicate that the filter is robust to small deviations in parameter values, and performance drops only under large perturbations.
Figure 3: Performance of different parameter combinations in terms of tracking distance and coverage. Colors represent distance from the analytically derived parameter set.
The narrow performance variance across trials and the existence of a stable parameter zone around the analytical parameter set indicate that the algorithm is stable under parameter perturbations.
All code, datasets, and evaluation scripts are publicly available at our git repository. This includes raw LiDAR point clouds, GPS ground truth, parameter configuration files, and plotting tools.
III-B Ablation study
In this subsection, we evaluate the individual components of our proposed algorithms to assess their contributions to overall performance. To isolate the effect of the moving LiDAR, we performed a control experiment where a static, non-moving LIVOX AVIA sensor was positioned next to the moving platform and recorded the same trajectory used in Subsection III-A. The number of returned points for both configurations is shown in FIGURE 4, where a significant difference can be observed, particularly in the furthest sections of the trajectory.
Figure 4: Returned point counts and trajectory for the moving and fixed platforms for the ’Horizontal fast’ trajectory (the left of this figure is the same as the far right of FIGURE 2). A notable difference is observed in the furthest sections of the trajectory, where the stationary LiDAR returns almost no points beyond 100m100\;m.
We compare the proposed delta position and constant velocity motion model PFs with the SOTA method of Dogru et al. [13] using an EKF with a constant velocity motion model. TABLE II presents the results of our ablation experiments. Each experiment was repeated 5 times to account for the non-deterministic nature of the methods.
The baseline maximum tracking distance of 70m70\;m is the reported result of [13], which was recorded using a circular rotating Velodyne VLP-16 LiDAR that tracked a drone with $s_{diag}=0.45\;m$, slightly larger than our DJI Phantom 4. When using the same EKF algorithm within the proposed tracking system, which utilizes the rosette scanning pattern of the LIVOX AVIA sensor, the tracking distance improves by approximately $15\;\%$, which we attribute to the sensor’s superior performance. Our particle filters extend this distance by an additional $30\;\%$ in the fixed LiDAR configuration.
In case of the moving platform, the maximum tracking distances are increased by another $30\;\%$, and are comparable across the EKF and both PFs. However, tracking coverage with the PFs is more than five times higher. We attribute the additional $30\;\%$ performance gain in maximum tracking distance to our strategy of moving the platform to maintain the drone within the center of the FoV. Overall, the best combination of maximum tracking distance and track coverage is achieved by the delta position PF.
TABLE II: Results of the ablation study.
Method
Coverage $\%$]
mean / std
Max Dist. $[m]$
mean / std
Fix, EKF [13]
reported with VLP-16
$N.A.$ $\sim 70/N.A.$
Fix, EKF [13]
$6.60/0.03$ $80.3/0.0$
Fix, PF constant velocity
$43.91/1.37$ $103.79/0.88$
Fix, PF delta position
$51.59/4.47$ $102.83/0.70$
Moving, EKF [13]
$12.70/0.42$ $131.70/5.33$
Moving, PF constant velocity
$68.42/1.00$ $129.16/0.38$
Moving, PF delta position
$71.35/6.07$ $129.68/0.36$
III-C Tracking performance comparision
The previously tuned particle filters, utilizing constant velocity and delta position motion models, were benchmarked against the EKF tracking methods presented by Dogru et al. [13] and Catalano et al. [24]. These comparisons were conducted within the proposed rosette scanning pattern LiDAR system, where the turret was controlled using GT GPS positions streamed from the drone, ensuring the drone remained centered within the FoV. Four different trajectories were considered and recorded on a sunny day. These are shown in FIGURE 2, along with the corresponding returned point counts.
TABLE III summarizes the results of this set of experiments. Each test was executed 5 times to account for the non-deterministic nature of the filters. We report both the mean and standard deviation for maximum tracking distance and track coverage metrics.
TABLE III: Maximum detection distance and track coverage results on four outside trajectories. Bold font denotes the best result in its column. Our PF using the delta position model performs best in terms of coverage in all scenarios, while it matches the SOTA in maximum tracking distance.
Simple distance Vertical Horizontal slow Horizontal fast
Method
Coverage $\%$]
mean / std
Max Dist. $[m]$
mean / std
Coverage $\%$]
mean / std
Max Dist. $[m]$
mean / std
Coverage $\%$]
mean / std
Max Dist. $[m]$
mean / std
Coverage $\%$]
mean / std
Max Dist. $[m]$
mean / std
EKF [24]
2.60 / 0.06 14.09 / 0.00 4.41 / 0.21 16.65 / 0.09 2.61 / 0.03 10.88 / 0.32 0.00 / 0.00 0.00 / 0.00
EKF [24] mod.
9.83 / 0.21 32.51 / 0.40 20.38 / 0.11 59.42 / 0.55 31.09 / 0.01 71.44 / 0.00 15.20 / 0.00 44.75 / 0.00
EKF [13]
37.80 / 1.56 131.23 / 0.00 7.55 / 0.03 51.21 / 0.00 14.08 / 0.03 138.18 / 0.00 12.70 / 0.42 131.71 / 5.33
PF constant
velocity
43.02 / 0.84 131.23 / 0.00 55.84 / 0.51 140.18 / 9.15 73.05 / 0.63 132.02 / 2.28 68.42 / 1.00 129.16 / 0.38
PF delta position 55.19 / 7.53 131.97 / 2.02 80.35 / 5.25 129.93 / 0.28 74.72 / 3.20 135.57 / 1.61 71.35 / 6.07 129.68 / 0.36
The EKF in [24] was originally tuned for indoor scenarios with static LiDAR placement, where the drones operate at lower speeds. Their method employs a Nearest-Neighbor Search (NNS) algorithm for finding the next possible drone measurement with a fixed radius ($r=s_{diag}/2=0.2\;m$) to associate measurements. We benchmarked our algorithm against their original method (EKF [24] in TABLE III) and a modified one, where we increased the NNS search radius (EKF [24] mod. in TABLE III). It was increased to r=1.5mr=1.5\;m, since this is the expected maximum travel distance based on the drone dynamics in one iteration of the filter. The EKF filters were all re-initialized at a specific time, when the LiDAR measurement only contained points from the drone. This adjustment was made because their original papers used different approaches for the initial detection, so to ensure a fair comparison, a proper initialization was provided. If the drone was successfully tracked for at least five consecutive frames, results are included in the table, otherwise, a value of 0.00 is reported.
The results indicate that our PF-based algorithms achieve consistently better track coverage than the EKF-based methods. The EKF introduced in [24] with the increased NNS radius performs well at close ranges. However, once the drone exits the expected radius, the tracker fails and cannot recover. [13] partially solves this problem by dynamically setting the NNS radius, which allows their EKF to occasionally reacquire the drone. Nevertheless, after that, their EKF prediction is usually inaccurate, and the drone is lost again in a few frames. This is reflected in the lower coverage values shown in TABLE III, except for the ’simple distance’ trajectory, where the motion is constant and predictable, enabling successful recovery. Our PF-based solutions adapt more effectively to the combination of sensor modality, drone dynamics, and missing measurements, thus having significantly better coverage across all scenarios.
Due to the unpredictable nature of drone motion, out of the two particle filters, the delta position-based solution performs consistently better in terms of coverage. It is somewhat interesting that it also outperforms the model-based one on the ’simple distance’ trajectory, where the speed is almost constant. This is likely because the high velocity noise (which is needed to track the drone when it changes direction quickly) results in bigger position errors, ultimately reducing its coverage.
Overall, we found that our particle filter with delta position motion model matches the SOTA EKF-based solution in terms of maximum tracking distance, while providing significantly better coverage. This is particularly important in real-world scenarios, where the turret is moved based on the detection of the filter to keep the drone in the middle of the FoV.
III-D Field test
We performed two measurements with the presented delta position PF method under two different weather conditions (on a sunny and a foggy day). The purpose of the validation is to test the track movement and to determine the detection distances and number of detected points in real-world conditions with turret input from the filter.
The same DJI Phantom 4 was used for these experiments. The measurements were carried out in the factory state of the drone, we did not improve the detection range by using special paint or other methods. For a video of the outdoor experiments, kindly check the video attachment here: https://youtu.be/ghfjqAnDuag.
A sunny day was chosen for the first and a foggy day for the second measurement day, when visibility was also severely limited (see FIGURE 5).
(a) Sunny
(b) Foggy
The drone’s trajectory was successfully tracked up to a distance of 128m128\;m in the sunny test environment (FIGURE 6(a)), and up to a distance of 58m58\;m in the foggy test environment (FIGURE 6(b)). As explained earlier in subsection II-C1, we consider the drone tracked when $\sigma_{particles}<0.3\;m$. It is important to note that we had some occasional detections at distances of up to 145m145\;m on the sunny day.
In both scenarios, we calculated maximum speeds of a little more than 6m/s6\;m/s from the tracking positions and timestamps.
Another important aspect is the number of detected points, which greatly impacts the robustness of tracking. In sunny weather, except for the last part of the detection range, the number of detected points in the result is $3$ to $5$ times higher than in the current SOTA (see FIGURE 7). This can be attributed partly to the sensor choice and partly to our strategy of maintaining the drone within the center of the FoV based on our findings in Subsection III-B.
Figure 7: Number of detected points depending on distance. The proposed algorithm returns considerably more points in every distance range than Dogru et al.[13].
In the case of foggy weather, the detection distance was significantly reduced, as expected in advance based on the literature[41]. While a detailed investigation of the performance in different weather conditions is beyond the scope of this work, we include this observation as an incidental but noteworthy outcome of our study.
III-E Accuracy
These experiments were conducted within the MIMO arena [26] within the HUN-REN SZTAKI Institute for Computer Science and Control incorporating an OptiTrack motion capture system in a $7\times 8\;m^{2}$ office space with a height of 3.5m3.5\;m, which can provide sub-millimeter accurate, 240 Hz tracking data for bodies with IR retroreflective markers. For the target of the experiment a Crazyflie 2.0 MAV by Bitcraze was used, which measures $s_{diag}=12\;cm$. The expected maximum velocity and acceleration were 1.2m/s1.2\;m/s and $1.2\;m/s^{2}$ respectively, so the following noise parameters were set for the PF: $\sigma_{pos}=0.04\;m$, $\sigma_{vel}=0.04\;m$, $\sigma_{m}=0.06\;m$.
@@ III-E Accuracy
Figure 8: Initial detection error and $\sigma_{particles}$ of the particle filter. The particle filter locks onto the drone almost instantaneously after receiving the first measurement at around 2 seconds. $\sigma_{threshold}=\sigma_{m}\cdot 1.5=0.09\;m$ is also plotted.
@@
FIGURE 8 shows the moment the drone gets tracked by the delta position PF, by plotting the error as the distance of the estimated state to the true state in meters of a representative detection. The average of the standard deviation of the particles in each dimension is also plotted (denoted as $\sigma_{particles}$).
@@
While the drone is out of the FoV of the LiDAR sensor, the particles remain idle, only influenced by the added noise in the predict step which results in a slowly inflating particle cloud with a slowly rising $\sigma_{particles}$. In this case, the MAV is resting on the floor and, because of its small size, is hidden inside the background model. As the drone starts hovering and measurements of it arrive at around 2s2\;s the particles gather and allow for the first good estimation. Parallel with this, the $\sigma_{particles}$ of the estimation drops sharply. This means the drone is tracked.
@@
For calculating the accuracy, the MAV was flown following certain flight patterns (described in FIGURE 9) in a repeated manner. The ground truth data was obtained by the motion capture system within the MIMO arena by recording the pose of the LIVOX sensor and the position of the drone and aligning the LIVOX measurements with the MIMO arena using the recorded pose. The patterns in FIGURE 9 were drawn from the point of view of the LiDAR sensor where the middle of each trajectory in the horizontal direction was in front of the sensor with a minimum distance of around 3m3\;m. Each circle represents a stationary wait time of 2s2\;s and the one with a black outline represents the first position of the respective pattern. The trajectory represented by the arrows in-between was flown in 3s3\;s in case of the vertical, horizontal, and ’lost and found’ pattern which resulted in a maximum speed of about 0.6m/s0.6\;m/s, while it was flown in 2.25s2.25\;s in case of the fast pattern which resulted in a maximum speed of around 1.2m/s1.2\;m/s. In each experiment, the drone flies to the first point and then it flies the pattern 3 times in the case of the vertical and horizontal patterns and 4 times in the case of the others.
TABLE IV: Accuracy comparison for different tracking methods. RMSE and 95 % confidence intervals are reported across five runs for each trajectory. Our delta position PF consistently achieves the lowest errors, which is particularly apparent in the ‘Fast’ and ‘Lost and found’ scenarios.
Horizontal Vertical Fast Lost and found
Method RMSE $[m]$
CI 95% $[m]$
lower / upper
RMSE $[m]$
CI 95% $[m]$
lower / upper
RMSE $[m]$
CI 95% $[m]$
lower / upper
RMSE $[m]$
CI 95% $[m]$
lower / upper
EKF [24]
0.078 0.076 / 0.081 0.595 0.577 / 0.614 2.604 2.474 / 2.749 1.880 1.810 / 1.955
EKF [13]
0.072 0.070 / 0.075 0.071 0.069 / 0.073 0.239 0.227 / 0.252 0.800 0.771 / 0.833
PF constant
velocity
0.080 0.078 / 0.082 0.077 0.075 / 0.080 0.156 0.147 / 0.166 0.643 0.619 / 0.669
PF delta position 0.054 0.053 / 0.056 0.057 0.055 / 0.059 0.132 0.125 / 0.139 0.292 0.281 / 0.304
TABLE IV presents a comparison of the accuracy results for the different tracking filters. As before, all tests were conducted five times to account for the non-deterministic nature of the filters. To ensure a fair comparison, all filters were properly initialized. We report the Root Mean Square Error (RMSE) over all five runs, calculated using the Euclidean distance as defined in (13) where $(\hat{x}_{i},\hat{y}_{i},\hat{z}_{i})$ denotes the the filter’s prediction, $(x_{i},y_{i},z_{i})$ denotes the GT position, and $n$ is the number of GT measurements. The $95\;\%$ confidence intervals were computed analytically using the Chi-Squared method. The ’Lost and found’ trajectory is discussed separately in Subsection III-F.
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(\sqrt{(\hat{x}_{i}-x_{i})^{2}+(\hat{y}_{i}-y_{i})^{2}+(\hat{z}_{i}-z_{i})^{2}}\right)^{2}}$ (13)
In these experiments, we also evaluate the proposed method against the EKF implementation from Catalano et al.[24] and the EKF approach proposed by Dogru et al.[13]. In the case of [24], we used their original NNS search radius, which was tuned for indoor scenarios and did not modify it for this evaluation.
For the ’Horizontal’ trajectory, the observed results are consistent with those reported in the respective papers. However, in the ’Vertical’ and ’Fast’ trajectories, the EKF in [24] performs considerably worse than the other methods. This poor performance is likely due to the motion model they employed: a Constant Turn Rate and Velocity (CTRV) model extended with a constant vertical velocity component equal to that of the planar motion. This results in degraded estimation along the vertical axis. In the ’Fast’ trajectory, tracking is lost almost immediately, which accounts for the high RMSE values there.
The proposed delta position PF consistently outperforms all other methods across all evaluated scenarios. This demonstrates that our approach not only matches the SOTA in terms of maximum tracking distance and surpasses it significantly in terms of track coverage, but is also the most accurate overall.
@@
In FIGURE 10, the Euclidean error of the delta position particle filter (first of the 5 runs in every scenario) is plotted with the ground truth speed calculated from the MIMO measurements. It is visible that the tracking is stable over time with an average position error of around 5cm5\;cm in the case of the horizontal and vertical measurements, while it is around 11cm11\;cm for the fast pattern.
@@
Figure 10: Position error of the particle filter plotted together with ground truth speed from the MIMO arena for the Vertical (top), Horizontal (middle), and Fast (bottom) patterns. The mean ($\mu_{error}$) and standard deviation ($\pm\sigma_{error}$) of the particle filter error are also plotted.
@@
It can be seen on all the plots of FIGURE 10, that the error of the particle filter is dependent on the speed. When the drone is at a waypoint, not moving, the error of the filter is around $3$ to 5cm5\;cm, while at the peak speeds of 1.2m/s1.2\;m/s of the fast pattern (not counting the one outlier), the error is around 20cm20\;cm. This dynamic difference of about 15cm15\;cm in the error can mostly be explained by the different delays in the pipeline. The internal processing time of the LiDAR firmware is 2ms2\;ms based on the data-sheet, the LiDAR scan integration time is 100ms100\;ms, and there is some additional overhead in the formatting, stamping, and publishing of the data to ROS topics. The LiDAR callback took 1.45ms1.45\;ms on avarage while the particle filter (predict, update, and resample steps including the background subtraction) took 10.85ms10.85\;ms. The whole delay is about 120ms120\;ms, in which the drone travels roughly 14.4cm14.4\;cm, which accounts for most of the dynamic error.
It is important to note that the size of the Crazyflie drone ($s_{diag}=12\;cm$) is on the same order of magnitude as the observed error. The LiDAR only returns points from the side of the drone facing the sensor, and the markers used by the OptiTrack system for tracking are intentionally placed asymmetrically. As a result, the true center of the drone does not exactly coincide with the measured GT position. Although these factors are difficult to quantify precisely, they all contribute to the overall error.
@@
Overall, when the drone is stationary, the error remains small and consistent. However, during movement, a dynamic error arises that increases linearly with speed. This affects the tracking accuracy but does not compromise its robustness.
III-F Losing and regaining Tracking